OpenStudy (anonymous):

Which ordered pair is a solution of the equation 2x - 3y = 18?

OpenStudy (anonymous):

substitute 0 for x when solving for y then substitute 0 in for y when solving for x put in (x,y)

OpenStudy (gorv):

well u have to plugg in values and try

OpenStudy (gorv):

u can put any value of x and calculate value of y for that value of x

OpenStudy (gorv):

like x=2

OpenStudy (gorv):

2*2-3y=18 4-3y=18 subtract 4 from both side

OpenStudy (anonymous):

3 can not be divided into 14 with these questions you substitute 0 in as a value and solve for the other variable

OpenStudy (gorv):

well 14/3 is also a point

OpenStudy (gorv):

like 1/2=0.5 decimal number u heard of it

OpenStudy (anonymous):

they're asking for the ordered pair solution 3(0)-6y=18 -6y=18 y=-3 3x-4(0)=18 3x=18 x=6 ordered pair would be (6,-3) (This is an example not the actual problem)

OpenStudy (anonymous):

You don't know what x or y are at this point so you substitute in 0

OpenStudy (gorv):

it could be any thing (2,14/3 ) are also a pair

OpenStudy (gorv):

we can put any value of x here

OpenStudy (anonymous):

No you can't because you don't know the value. This needs to be substituted with 0 to find the values.

OpenStudy (anonymous):

@Tyeisha_baby is this multiple choice

OpenStudy (gorv):

we put x=2 first as u r putting x=0

OpenStudy (gorv):

thn we will get y value loll

OpenStudy (anonymous):

To get value of y we substitute in x as 0 ex 3(0)+6y=18 6y=18 The value of y will be 3

OpenStudy (anonymous):

to get the value of x we substitute in 0as y ex 3x+6(0)=18 3x=18 the value of x is 6